# S-box Security Analysis of NIST Lightweight Cryptography Candidates: A Critical Empirical Study

### v1

@@ 5 Conclusion
In this paper, we have thoroughly discussed $25$ plus well known properties of cryptographic sboxes and described their behaviour under affine transformations. In addition, we have also provided an insight into the minimum and maximum achievable bounds of these properties and the desirable bounds for an optimal sbox.
Our aim was to provide a comprehensive walk-through on the taxonomy of cryptographic properties of sboxes for researchers who are new to the field of design and cryptanalysis of symmetric and lightweight ciphers.
Since the initiation of NIST lightweight cryptography competition, many attacks have been published on different ciphers in the competition. Many of these attacks exploit one of the weakness in the substitution layer. We hope that our analysis of these sboxes and overview on the bounds, will serve as a pre-requisite profile for cryptanalysts to determine new attack distinguishers and propose countermeasures to make the ciphers secure against these attacks.

### v3

@@ 7 Conclusion
Since the announcement of NIST lightweight cryptography competition, a considerable number of attacks have been published for exploiting the vulnerabilities in the non-linear component of the competitors under consideration. This paper presents a detailed cryptographic security analysis of the S-boxes of six finalists, examining the $24$ cryptographic properties that are relevant to established attack methods (see3). These six S-boxes exhibit characteristics such as high SAC and BIC values, degree-2 polynomials, and undisturbed bits — that may increase the susceptibility to known differential, linear, and boomerang attacks. The existence of degree 2 polynomials is the primary source of linear structures, followed by undisturbed bits. Improving the S-box design may, therefore, help improve their overall security margin of these offered by these permutations.
8 Declarations
Author Contribution Statement:
All authors contributed equally to the writing of the manuscript.
Author Competing Interest statement: Sajid Ali Khan has position on editorial board of Discover Computing, and was not involved in the review or decisions related to this manuscript. The other authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Funding Declaration:
Not applicable
Ethics and Consent to Participate declarations:
Not applicable
Consent for publication Declaration:
Not applicable
Ethics, Consent to Participate, and Consent to Publish declarations: Not Applicable
Availability of Data: The data that support the findings of this study are available from the corresponding author, upon request.
Preprint: A preprint of this article is available at [131]